OpenStudy (anonymous):

Prove the statement using epsilon and delta definition of a limit lim as x -> -2 (x^2 -1) =3

OpenStudy (anonymous):

are u typing or is it a mistake?

OpenStudy (anonymous):

okay =)

OpenStudy (ybarrap):

Let epsilon > 0 Need to find a delta > 0 such that if |x - (-2)| < delta then | (x^2 - 1) - 3| < epsilon |(x^2 - 1 )- 3| = |x^2 - 4| = |(x + 2)(x - 2)| < epsilon means -epsilon < (x + 2)(x - 2)< epsilon (-epsilon + 4)/(x - 2) < x + 2 < (epsilon + 4)/(x - 2) -(epsilon + 4) < (-epsilon + 4)/(x - 2) < x + 2 < (epsilon + 4)/(x - 2) < epsilon + 4 ==> |x + 2| < epsilon + 4 so if delta = epsilon + 4, then |x^2 - 4| < epsilon Ans. Since delta exists, then lim x^2 - 1 = 3 as x -> -2

OpenStudy (anonymous):

can u plz help me w/ a limit problem? these are soo confusing!

OpenStudy (anonymous):

its similar to this

OpenStudy (ybarrap):

ok

OpenStudy (ybarrap):

Trick is to focus on the antecedent (i.e. the second part of the if then part) and simplify until you can get the precedent (i.e. the first part of the if then statement)

OpenStudy (ybarrap):

correction in the "If p then q", p is the antecedent and q is the consequent. Focus on q to get to p.

OpenStudy (ybarrap):

anyway, what's the problem

OpenStudy (anonymous):

We were learning about continuity. If f and g are continous functions with f(3) = 5 and lim as x-> 3 [2 f(x) - g(x)] =4, find g(3).

OpenStudy (ybarrap):

lim 2f(x) - g(x) = 2*lim f(x) - lim g(x) = 2* f(3) - lim g(x) = 2* 5 - lim g(x) = 4 2*5 - 4 = lim g(x) as x -> 3 lim g(x) = 10 -4 = 6 as x -> 3 Since lim g(x) as x->3 exists and g is continuous, this implies g(3) = lim g(x) as x->3 then g(3) = 6.

OpenStudy (anonymous):

oh i see...tahnks alot for your help. it made it much more clearer. =)